If Y varies inversely as X^2 and Y=2 X=4.
What is the constant of variation?

Answer by venugopalramana(3286)

Y VARIES INVERSELY AS X MEANS Y=K/X....WHERE K IS A CONSTANT...WE ARE GIVEN THAT Y=2...WHEN.X=4...SO
2=K/4...OR...K=2*4=8.........SO 8 IS THE CONSTANT OF VARIATION AND
Y=8/X